OpenStudy (anonymous):

factor: 6x^2 - 5x + 1

OpenStudy (anonymous):

(3x-1)(2x-1)=0 3x-1=0 x=(1/3) 2x-1=0 x=(1/2) So, if you need the x values they are (1/3) and (1/2)

OpenStudy (precal):



OpenStudy (precal):

I used a method called slide and divide